# Shadows of Loop Quantum Black Holes: Semi-analytical Simulations of Loop Quantum Gravity Effects on Sagittarius $*{}^{*}$ and M $*{}^{*}$

### 07 Dec 2023 v1

@@ 1 Introduction
One of the most effective ways to address these anomalies is to propose various classical modified theories of gravity or quantum gravity. Therefore, it is essential to test the modified theories beyond GR to confirm the final theory of gravity. The loop quantum gravity (LQG) elegantly resolves both the classical big bang and black hole singularities. Recently, a self-dual, regular, and static spacetime metric is obtained in the mini-superspace approach based on the polymerization procedure in LQG [10]. This self-dual spacetime is regular and free of any spacetime curvature singularity. In this spacetime, the effects of LQG are characterized by two parameters, i.e., the minimal area and the Barbero-Immirzi parameter. It is also shown that this spacetime is self-dual in the sense of T-duality [see 11, 12, for detail]. Over the past ten years or so, many studies have been done on black holes in LQG using various quantization methods. These black holes have been found to have excellent properties and not have any singularities at the origin [13, 14, 15, 16, 17, 18, 20, 19, 21, 22].
@@
A key question has emerged regarding whether the LQG effects in the loop quantum spacetime can produce any observational signatures for current and future experiments, allowing us to test the LQG directly. This has spurred a surge of studies over the past decades, drawing from various experiments and observations [e.g., 23, 24, 25, 28, 27, 29, 30, 31, 32, 33, 34, 12, 35, 26, 37, 36]. In [34], the LQG effects on the shadow of the rotating black hole are discussed in detail. They also subsequently discussed the observational implications of their results in the latest Event Horizon Telescope (EHT) observation of the supermassive black hole M87. Moreover, the gravitational lensing in the self-dual spacetime is studied, and the polymeric function from LQG is constrained by using the Very-Long-Baseline Interferometry (VLBI) data of the solar gravitational deflection of radio waves [12]. Recently, the solar system test and celestial mechanics test by S2 orbit around Sgr $*{}^{*}$ of the self-dual spacetime are considered in [35] and [38] respectively. They derived the observational constraints of the polymeric function $P$ of LQG. It is interesting to note that the phenomenological studies of other types of loop quantum black holes have also been widely explored [39, 40, 41, 42, 43]. However, there are still very few studies performed to extract the observable effects of rotating LQBH in a realistic astrophysical environment.
@@
Recently, the EHT collaboration released the first direct image of a black hole in human history, which captured a horizon scale image of a supermassive black hole at the center of a neighboring elliptical galaxy M87 [58, 59, 63, 61, 62, 60]. They found that the diameter of the center black hole shadow is $(42\pm 3)\;{\rm\mu as}$ with a deviation $\lesssim 10\%$ from circularity from this image. This leads to a measurement of the center mass, $M=(6.5\pm 0.7)\times 10^{9}M_{\odot}$ [58]. These results are in good agreement with the predictions of GR by assuming the geometry of black holes is described by the Kerr metric. Shortly thereafter, the EHT did horizon scale observation of Sgr $*{}^{*}$, which is the highest angular resolution observation so far [64, 65, 66, 67, 68, 69]. Sgr $*{}^{*}$ shows complicated features that remain unsettled and yet active subjects of study. The rapid intrinsic variation puts a great challenge to do a detailed observation of it [67]. Flares are observed at near-infrared and X-ray [e.g., 70, 71], which is also one of the major frontiers of investigation. Multiple astrophysical models have been proposed to explain the observed features of Sgr $*{}^{*}$ [72, 73, 74, 75]. All these studies, though, are predicated on the idea that the Kerr metric provides the geometry of the black hole. There has not been a significant deviation from GR detected yet in the constraints of EHT observation [69]. Nevertheless, the alternative theory of gravity cannot be ruled out [78, 79], and more study is required. The observations by EHT can still be used to probe the environment around the black hole and distinguish or constrain black holes in different theories of gravity [76, 77, 78, 80, 81, 63, 69].
@@ 2 A Brief Introduction to LQBH spacetime
Finally, the Immirzi parameter $(\gamma)$ has a lot of choices from different considerations [see 85, 86, 87, 88, 89, 90]. It has been shown that its value can even be a complex number [86, 85, 89, 91], or considered as a scalar field in which the value would be fixed by the dynamics [90]. We adopt the commonly used value $\gamma=0.2375$ from the black hole entropy calculation [91].
@@ 4.2 Degeneracy of blurred shadow
To understand more detail of the shadow image of Sgr $*{}^{*}$ with LQBH. We select 3 points on the $d_{\rm ring}=51.8\,\mu as$ contour of the blurred cases with an inclination angle of $30^{\circ}$, which correspond to similar ring diameters in the resolution of EHT (upper right panel of Fig. 5). Their parameters are $(a,P)=(0.69,0.00)$, $(0.66,0.08)$, and $(0.55,0.16)$, where the first number in the brackets is the value of BH spin while the second one is the value of polymeric function. The GRRT results of these parameters are presented in Fig. 6. The upper panels are the unblurred shadow images, while the lower ones are blurred results. As the $P$ increases, we can clearly see a larger photon ring. The diameters of the lensing ring are $51.45,54,25$ and $60.66\,\rm\mu as$. However, it is not clearly reflected in the corresponding blurred images. The radii of the outer horizons of the 3 black holes are $1.73\,\rm r_{\rm g}$, $1.79\,\rm r_{\rm g}$, and $1.87\,\rm r_{\rm g}$, corresponding to ring diameters of $51.45,51.05$, $51.05\,\rm\mu as$. To distinguish the degeneracy of spin and polymeric function, observations in higher resolution are required, which may be obtained from future EHT and ngEHT observations [102, 103]. From Fig. 6, we also see that the BH images in the right column are slightly dimmer than those of the other two columns. Comparing the blurred images, we find that the brightest pixel on the right panel ($a=0.55$, $P=0.16$, LQBH case) is about $77\%$ of the left panel ($a=0.69$, $P=0$, Kerr BH case).
The value of $\angle\beta_{2}$ is another channel to distinguish the Black Holes. The $\angle\beta_{2}$ that calculated from unblurred GRRT result are $125^{\circ}$, $123^{\circ}$, and $124^{\circ}$, and from the blurred result, $\angle\beta_{2}$ are $129^{\circ}$, $126^{\circ}$, and $123^{\circ}$. The difference of $\angle\beta_{2}$ is too small to be distinguished in current polarimetric observation [97]. Therefore the measurement of $\beta_{2}$ in the case of Sgr $*{}^{*}$ may not be enough to determine the $a$ and $P$ parameters. Observations with higher angular resolutions of Sgr $*{}^{*}$ will be important to find a more consistent upper limit of the LQG parameter $P$ in future surveys (e.g., ngEHT).
@@ 4.3 EHT constraint on the polymetric function of M $*{}^{*}$
It is indicated that the information of $\beta_{2}$ is important to determine the possible parameters for M $*{}^{*}$ and Sgr $*{}^{*}$. With the future polarimetric data of Sgr $*{}^{*}$, we will be able to make a more accurate constraint on the LQG effect of it. In addition to these constraints, the measurement of jet power can offer an additional robust constraint on the parameters of the black hole [62, 68]. Calculation of jet power requires explicit GRMHD simulations, which can directly compare the jet power from different BHs and compare it with the EHT constraint. We plan to do such studies in the future.
@@ 5 Summary and discussion
    By the current resolution of EHT, the constraint has some degeneracy between spin $a$ and polymetric function $P$. We cannot distinguish Kerr BH and LQBH only through the ring diameter. Future higher-resolution surveys (e.g., ngEHT) will provide stronger constrain on the LQG parameter.
@@
Finally, we would like to discuss some caveats of our study. Our study only uses a semi-analytical GRRT simulation. The higher-order corrections from the accretion flows is not included (e.g., hot spot, light curve variation, etc.), which we plan to explore in future studies with self-consistent 2D and 3D GRMHD simulations and GRRT calculations. On top of this, studying self-consistent spectral (e.g., image, spectra) and timing properties (e.g., variability) of astrophysical sources demands a two-temperature accretion flow model and different magnetic field configurations (e.g., [72, 110, 111]. We will carry forward such study in the future.
@@ Appendix A Geodesic motion in LQBH
$R(r)=0,\qquad\frac{dR(r)}{dr}=0.$ (A.12)

### 28 Dec 2023 v2

@@ 1 Introduction
One of the most effective ways to address these anomalies is to propose various classical modified theories of gravity or quantum gravity. Therefore, it is essential to test the modified theories beyond GR to confirm the final theory of gravity. The loop quantum gravity (LQG) elegantly resolves both the classical Big Bang and black hole singularities. Recently, a self-dual, regular, and static spacetime metric is obtained in the mini-superspace approach based on the polymerization procedure in LQG [9, 11]. This self-dual spacetime is regular and free of any spacetime curvature singularity. In this spacetime, the effects of LQG are characterized by two parameters, i.e., the minimal area and the Barbero-Immirzi parameter. It is also shown that this spacetime is self-dual in the sense of T-duality [see 10, 12, for detail]. Over the past ten years or so, many studies have been done on black holes in LQG using various quantization methods. These black holes have been found to have excellent properties and not have any singularities at the origin [13, 14, 15, 16, 17, 18, 19, 21, 20, 22, 23].
@@
A key question has emerged regarding whether LQG effects in the loop quantum spacetime can produce any observational signatures for current and future experiments, allowing us to test the effect of LQG. This has spurred a surge of studies over the past decades, drawing from various experiments and observations [e.g., 24, 26, 27, 30, 29, 31, 32, 33, 34, 35, 36, 12, 37, 28, 39, 38, 25]. In [36], LQG effects on the shadow of the rotating black hole are discussed in detail. They also subsequently discussed the observational implications of their results in the latest Event Horizon Telescope (EHT) observation of the supermassive black hole M87. Moreover, the gravitational lensing in the self-dual spacetime is studied, and the polymeric function from LQG is constrained by using the Very-Long-Baseline Interferometry (VLBI) data of the solar gravitational deflection of radio waves [12]. Recently, the solar system test and celestial mechanics test by S2 orbit around Sgr $*{}^{*}$ of the self-dual spacetime are considered in [37] and [40] respectively. They derived the observational constraints of the polymeric function $P$ of LQG. It is interesting to note that the phenomenological studies of other types of loop quantum black holes have also been widely explored [41, 42, 43, 44, 45]. However, there are still very few studies performed to extract the observable effects of rotating LQBH in a realistic astrophysical environment.
@@
Recently, the EHT collaboration released the first direct image of a black hole in human history, which captured a horizon scale image of a supermassive black hole at the center of a neighboring elliptical galaxy M87 [60, 61, 65, 63, 64, 62]. They found that the diameter of the center black hole shadow is $(42\pm 3)\;{\rm\mu as}$ with a deviation $\lesssim 10\%$ from circularity from this image. This leads to a measurement of the center mass, $M=(6.5\pm 0.7)\times 10^{9}M_{\odot}$ [60]. These results are in good agreement with the predictions of GR by assuming the geometry of black holes is described by the Kerr metric. Shortly thereafter, the EHT did horizon scale observation of Sgr $*{}^{*}$, which is the highest angular resolution observation so far [66, 67, 68, 69, 70, 71]. Sgr $*{}^{*}$ shows complicated features that remain unsettled and yet active subjects of study. The rapid intrinsic variation puts a great challenge to do a detailed observation of it [69]. Flares are observed at near-infrared and X-ray [e.g., 72, 73], which is also one of the major frontiers of investigation. Multiple astrophysical models have been proposed to explain the observed features of Sgr $*{}^{*}$ [74, 75, 76, 77]. All these studies, though, are predicated on the idea that the Kerr metric provides the geometry of the black hole. There has not been a significant deviation from GR detected yet in the constraints of EHT observation [71]. Nevertheless, the alternative theory of gravity cannot be ruled out [82, 83], and more study is required. The observations by EHT can still be used to probe the environment around the black hole and distinguish or constrain black holes in different theories of gravity [78, 79, 80, 81, 82, 84, 85, 65, 71].
@@ 2 A Brief Introduction to LQBH spacetime
Finally, the Immirzi parameter $(\gamma)$ has a lot of choices from different considerations [see 89, 90, 91, 92, 93, 94]. It has been shown that its value can even be a complex number [90, 89, 93, 95], or considered as a scalar field in which the value would be fixed by the dynamics [94]. We adopt the commonly used value $\gamma=0.2375$ from the black hole entropy calculation [95].
@@ 4.2 Degeneracy of blurred shadow
To understand more detail of the shadow image of Sgr $*{}^{*}$ with LQBH. We select 3 points on the $d_{\rm ring}=51.8\,\mu as$ contour of the blurred cases with an inclination angle of $30^{\circ}$, which correspond to similar ring diameters in the resolution of EHT (upper right panel of Fig. 5). Their parameters are $(a,P)=(0.69,0.00)$, $(0.66,0.08)$, and $(0.55,0.16)$, where the first number in the brackets is the value of BH spin while the second one is the value of the polymeric function. The GRRT results of these parameters are presented in Fig. 6. The upper panels are the unblurred shadow images, while the lower ones are blurred results. As the $P$ increases, we can clearly see a larger photon ring. The diameters of the lensing ring are $51.45,54,25$ and $60.66\,\rm\mu as$. However, it is not clearly reflected in the corresponding blurred images. The radii of the outer horizons of the 3 black holes are $1.73\,\rm r_{\rm g}$, $1.79\,\rm r_{\rm g}$, and $1.87\,\rm r_{\rm g}$, corresponding to ring diameters of $51.45,51.05$, $51.05\,\rm\mu as$. To distinguish the degeneracy of spin and polymeric function, observations in higher resolution are required, which may be obtained from future EHT and ngEHT observations [106, 107]. From Fig. 6, we also see that the BH images in the right column are slightly dimmer than those of the other two columns. Comparing the blurred images, we find that the brightest pixel on the right panel ($a=0.55$, $P=0.16$, LQBH case) is about $77\%$ of the left panel ($a=0.69$, $P=0$, Kerr BH case).
The value of $\angle\beta_{2}$ is another channel to distinguish the Black Holes. The $\angle\beta_{2}$ that calculated from unblurred GRRT result are $125^{\circ}$, $123^{\circ}$, and $124^{\circ}$, and from the blurred result, $\angle\beta_{2}$ are $129^{\circ}$, $126^{\circ}$, and $123^{\circ}$. The difference of $\angle\beta_{2}$ is too small to be distinguished in current polarimetric observation [101]. Therefore the measurement of $\beta_{2}$ in the case of Sgr $*{}^{*}$ may not be enough to determine the $a$ and $P$ parameters. Observations with higher angular resolutions of Sgr $*{}^{*}$ will be important to find a more consistent upper limit of LQG parameter $P$ in future surveys (e.g., ngEHT).
@@ 4.3 EHT constraint on the polymetric function of M $*{}^{*}$
It is indicated that the information of $\beta_{2}$ is important to determine the possible parameters for M $*{}^{*}$ and Sgr $*{}^{*}$. With the future polarimetric data of Sgr $*{}^{*}$, we will be able to make a more accurate constraint on LQG effect of it. In addition to these constraints, the measurement of jet power can offer an additional robust constraint on the parameters of the black hole [64, 70]. Calculation of jet power requires explicit GRMHD simulations, which can directly compare the jet power from different BHs and compare it with the EHT constraint. We plan to do such studies in the future.
@@ 5 Summary and discussion
    By the current resolution of EHT, the constraint has some degeneracy between spin $a$ and polymetric function $P$. We cannot distinguish Kerr BH and LQBH only through the ring diameter. Future higher-resolution surveys (e.g., ngEHT) will provide stronger constrain on LQG parameter.
@@
Finally, we would like to discuss some caveats of our study. The model in this work is not unique within the framework of LQG. In the LQG theory, different quantization schemes lead to black hole solutions with different spacetime properties, e.g. [9, 23, 12, 13, 14, 15]. Our result in this work implies the property of the metric in [36], they used the Newman-Janis method to generate a rotating loop quantum spacetime from a spherically symmetric self-dual black hole solution proposed in [9], where the polymeric function $P$ characterizes the loop quantum effect of this spacetime. The self-dual spacetime is derived in the mini-superspace approach based on the polymerization procedure in LQG [9]. It can be verified that under the transformation $r\rightarrow a_{0}/r$, where $a_{0}$ is related to the minimal area gap of LQG, the metric remains invariant. This is achieved with suitable re-parameterization of other variables, thus satisfying the T-duality [9, 12, 13]. We compute the shadow of this rotating LQBH in conjunction with simulations and examine the geometrical properties and polarization information of this shadow. From observations of the shadows of M87 and Sgr $*{}^{*}$ we find no significant traces of the loop quantum effect. For both of the BHs, the Kerr solution is in the allowed parameter space under the current observation limit. Thus we give an upper limit of $P\lesssim 0.2$ for the effect, i.e. if such a self-dual loop quantum effect exists, it should not be larger than this value.
Our study only uses a semi-analytical GRRT simulation. The higher-order corrections from the accretion flows are not included (e.g., hot spot, light curve variation, etc.), which we plan to explore in future studies with self-consistent 2D and 3D GRMHD simulations and GRRT calculations. On top of this, studying self-consistent spectral (e.g., image, spectra) and timing properties (e.g., variability) of astrophysical sources demands a two-temperature accretion flow model and different magnetic field configurations [e.g., 74, 114, 115]. We will carry forward such study in the future.
@@ Appendix A Geodesic motion in LQBH
$R(r)=0,\qquad\frac{dR(r)}{dr}=0,\qquad\frac{d^{2}R(r)}{dr^{2}}>0$ (A.12)